OpenStudy (anonymous):

Solve the inequality and type the solution in interval notation. No spaces! 7 ≤ 2x - 1 ≤ 9

OpenStudy (anonymous):

\[7+1\le2x-1+1\le9+1\]

OpenStudy (anonymous):

\[8\le2x\le10\]

OpenStudy (anonymous):

i need the notations ?

OpenStudy (anonymous):

\[4\le x \le 5\]

OpenStudy (anonymous):

\[ \left[ 4,5 \right]\]